How many grams of aluminum dichromate are formed from the double displacement reaction between 107.5g of aluminum nitrate and excess sodium dichromate? The products are aluminum dichromate and sodium nitrate (show work)
1
Expert's answer
2020-04-18T07:00:08-0400

The reaction can be shown as following:

2Al(NO3)3 + 3Na₂Cr₂O₇ -->  Al2(Cr2O7)3 + 6NaNO3

The amounts of aluminum dichromate produced in the reaction are the same as the amounts of aluminum nitrate used during the reaction. As a result:

m(Al2(Cr2O7)3) = m(Al(NO3)3) × Mr(Al2(Cr2O7)3) / [2 × Mr(Al(NO3)3)]= 107.5 g × 702 g/mol / [2 × 213 g/mol] = 177.15 g


Answer: 177.15 g